OpenStudy (anonymous):

Find, in square cm, the area of an equilateral triangle of side 20 cm.

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

Equilateral triangles are 60-60-60 trainagles which can be split into two 60-90-30 triagnles. Using 1/2base(height) you can solve this. If you draw this up for area you will be working with the opposite line and the adjacent so you want to work with tangent. Our adjacent line is hald the base of a 60-60-60 triangle and is 10cm. \[\tan 60 = Opposite / Adjacent\] \[\tan 60 = Opposite / 10cm\] \[Opposite = 10 \tan 60\] Plugging this back into the triangle area formula gets us (keeping in mind we double for two smaller triangles): \[2(1/2)(10cm)(10 \tan 60)cm = 173.21 or 100\sqrt{3} since \tan 60 = \sqrt{3}/1\]

OpenStudy (anonymous):

Forgive my typos :)

OpenStudy (anonymous):

but its result is 100(3)1/3

OpenStudy (anonymous):

ohhhhhhhhhhhhh ok I understand now. Thanks

OpenStudy (anonymous):

100(3)1/3 is the same ur result?

OpenStudy (anonymous):

Written as raised powers it's \[100(3)^{1/2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So yes. \[100(3)^{1/2} = 173.2 =100\sqrt{3}\]

OpenStudy (anonymous):

Some teachers want you to have the irregular expresion in there instead of just the calculated answer to prove you did the work.

OpenStudy (anonymous):

And that answer would be in \[cm ^{?}\]

OpenStudy (anonymous):

\[cm^{2} \]

OpenStudy (anonymous):

ok thanks